OpenStudy (anonymous):

1/3+?/6=2/3 would it be 3

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

please help me

OpenStudy (anonymous):

sure find the common denominator of all of them like this \[\frac{2}{6} + \frac{?}{6} = \frac{4}{6}\] does this help?

OpenStudy (anonymous):

yeah it would be 2

OpenStudy (anonymous):

i think i got it 2 is the answer

OpenStudy (anonymous):

yes sorry for the late response 2 is the answer